TUGBOAT

Volume 11, Number 1 ^{/} April 1990

3 Addresses

General Delivery 5 From the President ^{/} Nelson H. F. Beebe 6 From the past President and Annual Meeting host ^{/} Bart Childs 7 Donald E. Knuth Scholarship 8 Coordination of non-English use of TEX ^{/} Michael Ferguson 9 Using TEX 3 in a multilingual environment — Some ideas ^{/} Peter Breitenlohner 56 Solution to crossword of TUGboat vol. 10, no. 3

Software 12 Erratum: The new versions of TEX and , TUGboat vol. 10, no. 3 13 Virtual Fonts: More Fun for Grand Wizards ^{/} Donald Knuth 24 Additional Hyphenation Patterns ^{/} Gerard Kuiken

Fonts 26 Typesetting modern Greek — An update ^{/} Yannis Haralambous

Graphics 26 Combining graphics with TEX on IBM PC-compatible systems and LaserJet printers ^{/} Lee S. Pickrell

Resources 32 Summary of resources available to TEX users ^{/} Barbara Beeton and Ron Whitney 36 A proto-TUG bibliography ^{/} Barbara Beeton 38 The Clarkson archive server — Update as of 26 January 1990 ^{/} Michael DeCorte

· IBM VM/CMS 45 IBM VM/CMS site report ^{/} Dean Guenther

Typesetting on Personal Computers 46 Une version complète de TEX du domaine public pour compatibles PC : les “deux disquettes GUT” ^{/} Nicolas Brouard

Macros 51 Unusual Paragraph Shapes ^{/} Victor Eijkhout 54 TEXT1 goes public domain ^{/} Dean Guenther 57 Showing-off math macros ^{/} Michael J. Wichura 62 How to avoid writing long records to TEX’s \write streams ^{/} Peter Breitenlohner

Tutorials 62 Forward references and the ultimate dirty trick ^{/} Lincoln Durst

69 Output routines: Examples and techniques. Part I: Introduction and examples ^{/} David Salomon

L^{A}TEX 85 The L^{A}TEX column ^{/} Jackie Damrau

86 Announcing two reports: SGML-L^{A}TEX and Journal style guidelines ^{/}

C. G. van der Laan

87 International L^{A}TEX is ready to use ^{/} Joachim Schrod 91 The new font family selection — User interface to standard L^{A}TEX ^{/}

Frank Mittelbach and Rainer Schöpf

98 A style option to adapt the standard L^{A}TEX document styles to A4 paper ^{/}

Nico Poppelier and Johannes Braams

103 Some macros to draw crosswords ^{/} B Hamilton Kelly

Abstracts 120 Deutsche Kurzfassungen der TUGboat-Artikel ^{/} Luzia Dietsche

News & Announcements 122 TEX, TUG, and Eastern Europe ^{/} Hubert Partl 123 Calendar 125 GUTenberg’90, 15–17 May 1990, Toulouse, France 126 Call for papers: SGML & TEX conference, 31 August 1990, Groningen,

The Netherlands

Late-Breaking News 127 Production notes ^{/} Barbara Beeton

TUG Business 128 Institutional members

Forms 131 TUG membership application

Advertisements 148 Index of advertisers

Volume 11, Number 2 ^{/} June 1990

151 Addresses

General Delivery 153 Editorial comments ^{/} Barbara Beeton

154 TEX in the production environment — questionnaire responses ^{/} Elizabeth Barnhart

Software 165 Exercises for TEX: The Program ^{/} Donald Knuth

Philology 171 Character set encoding ^{/} Nelson Beebe 175 On standards for computer modern font extensions ^{/} Janusz S. Bień

Fonts 183 Circular reasoning: typesetting on a circle, and related issues ^{/} Alan Hoenig

Graphics 190 On the implementation of graphics into TEX ^{/} Gerhard Berendt 194 Including Macintosh graphics in L^{A}TEX documents ^{/} Len Schwer 200 Combining graphics with TEX on PC systems with laser printers, part II ^{/}

Lee S. Pickrell

Resources · Data General 206 Data General site report ^{/} Bart Childs · IBM VM/CMS 207 IBM VM/CMS site report ^{/} Joachim Lammarsch

207 Resources available to TEX users ^{/} Barbara Beeton 208 A proto-TUG bibliography: Installment two ^{/} Barbara Beeton 210 New books on TEX ^{/} Victor Eijkhout 211 DECUS TEX collection — submissions wanted ^{/} Ted Nieland

Questions 212 Just plain Q&A ^{/} Alan Hoenig

Tutorials 212 Output routines: Examples and techniques. Part II: otr techniques ^{/} David Salomon

Macros 237 Lists in TEX’s mouth ^{/} Alan Jeffrey 245 A nestable verbatim mode ^{/} Philip Taylor 250 Easy Table ^{/} Khanh Ha 265 Typesetting bridge via TEX ^{/} Kees van der Laan

L^{A}TEX 276 The L^{A}TEX column ^{/} Jackie Damrau 279 Making 35mm colour slides with SliTEX ^{/} Ken Yap 280 An easy way to make slides with L^{A}TEX ^{/} Georg Denk 284 A new implementation of the L^{A}TEX verbatim and verbatim* environments ^{/}

Rainer Schöpf

297 Reprint: The new font family selection — User interface to standard L^{A}TEX ^{/} Frank Mittelbach and Rainer Schöpf

Abstracts 305 Deutsche Kurzfassungen der TUGboat-Artikel ^{/} Luzia Dietsche

News & Announcements 308 Calendar 309 Production notes ^{/} Barbara Beeton 310 New publications and software available through TUG 313 Russians visit TUG headquarters

TUG Business 314 Institutional members

Forms 317 TUG membership application

Advertisements M–10 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes M–1 TUG Membership List Supplement

Volume 11, Number 3 ^{/} September 1990 1990 Annual Meeting Proceedings

Introduction 335 Nelson Beebe ^{/} President’s introduction

TEX & TEX-Based Systems 337 Frank Mittelbach ^{/} E-TEX: Guidelines for future TEX 346 Hisato Hamano ^{/} Vertical typesetting with TEX 353 Yoshiyuki Miyabe, H. Ohta, K. Tsuga ^{/} Structured document preparation system

AutoLayouter

Macros & Other Tools 359 Amy Hendrickson ^{/} Getting TEXnical: Insights into TEX macro writing techniques 371 Stephen Alex Fulling ^{/} Where’s the Greek shift key? 373 Micah Beck and A. Siegel ^{/} TransFig: Portable graphics for TEX 381 Andrew Marc Greene ^{/} BASIX: An interpreter written in TEX

Book Production 393 Helen Gibson ^{/} A noddy’s guide to high quality TEX production 403 Robert A. Adams ^{/} Problems on the TEX/PostScript/graphics interface 409 Stephan von Bechtolsheim ^{/} TEX in practice: Comments on a 4-volume, 1400-page

series on TEX

413 Mimi L. LaFrenz ^{/} Textbook publishing — 1990 and beyond

Preprocessors & Editors 417 David Ness ^{/} Diagnosing TEX errors with a preprocessor

421 Linda Williams and Linda Hall ^{/} Increased efficiency using advanced EDT editing features

Teaching & Training 425 Charles R. Martin ^{/} TEX for TEXnical typists

Fonts 429 Michael Vulis ^{/} VTEX enhancements to the TEX language 435 Alan Hoenig ^{/} A constructed Dürer alphabet

Abstracts 439 Luzia Dietsche ^{/} Deutsche Kurzfassungen der TUGboat-Artikel

Status Reports 442 Nelson Beebe ^{/} TEX 3.0 and 2.0 444 Frank Mittelbach ^{/} L^{A}TEX 2.10

International Reports 444 Joachim Lammarsch ^{/} dante, Deutschsprachige Anwendervereinigung TEX e.V. 445 Bernard Gaulle ^{/} News from and about GUTenberg 446 Kees van der Laan ^{/} NTG’s second year 447 Roswitha Graham and Jan Michael Rynning ^{/} Report from the Nordic TEX Users

Group

448 Malcolm Clark ^{/} TEX in the UK 450 Malcolm Clark ^{/} TEX in Europe

Site Reports 451 Craig Platt ^{/} MVS site report 451 David Kellerman ^{/} VMS site report 452 Bart Childs ^{/} Data General site report 453 John Crawford ^{/} Prime 50 series site report 453 Pierre A. MacKay ^{/} UNIXTEX site report 454 Joachim Lammarsch ^{/} VM/CMS site report

Q & A 455 Barbara Beeton ^{/} Report from the question and answer session

Participants 458 List of participants at the Texas A&M meeting (June 17–20, 1990)

Announcements 462 Calendar

TUG Business 463 Institutional members

Forms 466 TUG membership application

Advertisements 479 Index of advertisers

Volume 11, Number 4 ^{/} November 1990

483 Addresses

General Delivery 485 From the President ^{/} Nelson H. F. Beebe 488 The future of TEX 489 The future of TEX and ^{/} Donald E. Knuth 490 Comments on the future of TEX and ^{/} Nelson H. F. Beebe 494 Editorial comments ^{/} Barbara Beeton 497 Arthur Lee Samuel, 1901–1990 ^{/} Donald Knuth

Software 499 Answers to Exercises for TEX: The Program ^{/} Donald Knuth 511 Webless literate programming ^{/} Jim Fox 513 A TEX previewer for “slow” terminals ^{/} Harold T. Stokes

Philology 514 Report on multilingual activities ^{/} Michael Ferguson

Fonts 517 Filenames for fonts ^{/} Karl Berry 520 Arabic, Persian and Ottoman TEX for Mac and PC ^{/} Yannis Haralambous 525 Environment for translating to PostScript ^{/}

Shimon Yanai and Daniel M. Berry

542 An improved chess font ^{/} David Tofsted

Output Devices 545 TEX output devices (with charts) ^{/} Don Hosek

570 Report from the DVI Driver Standards Committee ^{/} Don Hosek

Resources 570 Updates from all over ^{/} Barbara Beeton 571 eplain ^{/} Karl Berry 572 New Books on TEX ^{/} Victor Eijkhout 573 A proto-TUG bibliography: Installment three ^{/} Barbara Beeton

· VAX/VMS 576 The 1990 DECUS TEX collection ^{/} Ted Nieland

578 The IVRITEX Mailing List ^{/} Don Hosek

· IBM VM/CMS 578 IBM VM/CMS site report ^{/} Joachim Lammarsch Typesetting on PCs 578 TEX implementations for IBM PCs: comparative timings ^{/} Erich Neuwirth

Tutorials 580 Long-winded endnotes and exercises with hints or solutions ^{/} Lincoln Durst

588 Output routines: Examples and techniques. Part III: insertions ^{/} David Salomon

Macros 605 A new editor ^{/} Victor Eijkhout 605 Line breaking in \unhboxed text ^{/} Michael Downes 612 Looking ahead for a ⟨box⟩ ^{/} Sonja Maus 613 An indentation scheme ^{/} Victor Eijkhout 616 A \parskip scheme ^{/} Victor Eijkhout 620 Sanitizing control sequences under \write ^{/} Ron Whitney 623 An Overview of EDMAC: A plain TEX format for critical editions ^{/}

John Lavagnino and Dominik Wujastyk

L^{A}TEX 644 The L^{A}TEX column ^{/} Jackie Damrau

644 A L^{A}TEX document style option for typesetting APL ^{/} Andreas Geyer-Schulz,

Josef Matulka and Gustaf Neumann

652 Experiments in TEXnicolour — A SliTEX sub-style for Colour Printers ^{/} David Love 657 Footnotes in a multi-column layout ^{/} Frank Mittelbach

Abstracts 663 Deutsche Kurzfassungen der TUGboat-Artikel ^{/} Luzia Dietsche

Letters 665 TEX drug bug ^{/} Peter C. Akwai

News & Announcements 666 Calendar

cover 3 Call for papers: TUG Twelfth Annual Meeting, suburban Boston, 15–18 July 1991 667 Call for papers: Tenth meeting, DANTE, Vienna, 20–22 February 1991 667 TEX91/Congres GUTenberg’91, Paris, 23–26 September 1991 668 Call for papers: RIDT 91 (raster imaging/digital typography), Boston, 15–16 October 1991

TUG Business 669 TUG Bylaws

676 Institutional members

Late-Breaking News 679 Production notes ^{/} Barbara Beeton

679 Coming next issue

Forms 681 TUG membership application

Advertisements 693 Index of advertisers Supplements Computers & Typesetting: Errata, Supplement; TUG Membership List Supplement

Volume 12, Number 1 ^{/} March 1991 TEX90 Conference Proceedings

Introduction 3 Introduction ^{/} Barbara Beeton

Databases and Hypertext and their Relationship with TEX 5 Quick and dirty databases with nice output: AWK and TEX ^{/} Erich Neuwirth 8 TEX & hypertext — The future of electronic publishing? ^{/} Christine Detig 13 Experiments with TEX and hyperactivity ^{/} Les Carr, Sebastian Rahtz and Wendy Hall

User Groups and Dissemination of Information 21 The Dutch national L^{A}TEX effort ^{/} Johannes Braams, Victor Eijkhout

and Nico Poppelier

25 Documenting a TEX archive ^{/} Adrian Clark

Typographics: Æsthetics and Practicalities 28 L^{A}TEX-paragraphs floating around figures ^{/} Thomas Kneser 31 The document style designer as a separate entity ^{/} Victor Eijkhout 35 Improving the æsthetics of mixed-font documents ^{/} Philip Taylor 45 ArchıTEX, a preliminary international page pattern maker ^{/} Alan Wittbecker 58 Integration of graphics into TEX ^{/} Friedhelm Sowa

Expanding Horizons 64 PostScript, QuickDraw, TEX ^{/} Timothy Murphy

66 An approach to drawing circuit diagrams

for text books ^{/} M.P. Maclenan and G.M. Burns

70 A graphic driver to interface statistical software S and PICTEX ^{/} Olivier Nicole

Late Breaking News 74 Towards L^{A}TEX 3.0 ^{/} Frank Mittelbach and Rainer Schöpf

Public Domain TEX 80 Public-domain, documented implementations of TEX and for VAX/VMS ^{/} Brian Hamilton Kelly

Documentation, Origination, Editing and Markup 84 Post congress tristesse ^{/} Malcolm Clark 90 SGML (, TEX and …) ^{/} Kees van der Laan 105 SGML and TEX in scientific publishing ^{/} N.A.F.M. Poppelier 110 Getting TEXnical: Insights into TEX macro writing techniques ^{/} Amy Hendrickson 111 BibTEX reconsidered ^{/} Reinhard Wonneberger and Frank Mittelbach

Font Design and New Fonts 125 Labelling figures in TEX documents ^{/} Alan Hoenig

129 Typesetting old german: Fraktur, Schwabacher, Gotisch and initials ^{/}

Yannis Haralambous

139 The Irish alphabet ^{/} Mícheál Ó Searcóid 149 An international phonetic alphabet ^{/} Dean Guenther and Janene Winter 157 Practical halftoning with TEX ^{/} Adrian F. Clark

TEX Training and Education: The New User 166 Some TEX manuals ^{/} Angela Barden 171 TEX in schools: Just say no ^{/} Konrad Neuwirth

Announcements 175 Calendar 176 Nominations 176 TEX in Czechoslovakia ^{/} Jiří Veselý 177 sbTEX and sbMF ^{/} Wayne Sullivan 178 TEX TALK ^{/} T.V. Raman

TUG Business 179 Institutional members

Forms 182 TUG membership applications

Advertisements 200 Index of advertisers

Volume 12, Number 2 ^{/} June 1991

203 Addresses

General Delivery 205 President’s introduction ^{/} Nelson H. F. Beebe 208 Editorial comments ^{/} Barbara Beeton 211 TEX in Germany ^{/} Walter Obermiller

Philology 212 Russian TEX ^{/} Basil Malyshev, Alexander Samarin and Dimitri Vulis 215 Serbo-Croatian hyphenation: a TEX point of view ^{/} Cvetana Krstev 224 On TEX and Greek… ^{/} Yannis Haralambous 227 JemTEX 2.00 available for Japanese ^{/} François Jalbert

Fonts 227 Labelled diagrams in ^{/} Alan Jeffrey Graphics 229 X bitmaps in TEX ^{/} Reinhard Fößmeier

Output Devices 232 Report on the DVI Driver Standard ^{/} Joachim Schrod

Resources 233 Review of 3 : 16 Bible Texts Illuminated ^{/} Karl Berry and Kathy Hargreaves

235 Review of L^{A}TEX for engineers and scientists ^{/} Nico Poppelier

Questions 237 Just plain Q&A ^{/} Alan Hoenig Tutorials 238 The \if, \fx and \ifcat comparisons ^{/} David Salomon

248 Some tools for making indexes: Part I ^{/} Lincoln Durst 253 The structure of the TEX processor ^{/} Victor Eijkhout

Warnings 257 Initiation rites ^{/} Barbara Beeton

259 Solution to the riddle from TUGboat 11, no. 4 ^{/} Frank Mittelbach

Macros 260 The bag of tricks ^{/} Victor Eijkhout 261 TEX macros for producing multiple-choice tests ^{/} Don De Smet 270 Getting \answers in TEX ^{/} Jim Hefferon 272 Oral TEX ^{/} Victor Eijkhout 277 An expansion power lemma ^{/} Sonja Maus 278 Generating \n asterisks ^{/} George Russell 279 TEX and envelopes ^{/} Dimitri Vulis

L^{A}TEX 284 The L^{A}TEX column ^{/} Jackie Damrau 285 A comment on The L^{A}TEX column ^{/} Nico Poppelier 286 L^{A}TEX tree drawer ^{/} Glenn L. Swonk 290 “See also” indexing with Makeindex ^{/} Harold Thimbleby 291 Babel, a multilingual style-option system for use with L^{A}TEX’s standard document

styles ^{/} Johannes Braams

Queries 302 Public domain SGML tools wanted ^{/} Jeff Lankford

302 Reporting TEX’s hyphenations ^{/} Jiří Veselý

Letters 302 Response to Victor Eijkhout ^{/} Paul Abrahams 303 Response to Paul Abrahams ^{/} Victor Eijkhout 303 TEX in schools: Why not? ^{/} Al Cuoco

Abstracts 305 Cahiers GUTenberg #7 and #8 News & Announcements 307 Calendar

309 TEX91/Congres GUTenberg’91, Paris, 23–26 September 1991 311 Desktop Publishing in astronomy and space sciences, Strasbourg, 1–3 October 1991 311 Call for papers: EP92: International conference on electronic publishing, document

manipulation, and typography, Lausanne, Switzerland, 7–10 April 1992

Late-Breaking News 313 Fixed-point glue setting: Errata ^{/} Donald Knuth 313 Production notes ^{/} Barbara Beeton 315 Coming next issue

TUG Business 315 TUG financial statements 319 Institutional members 321 TUG Bylaws

Forms 331 TUG membership application

Advertisements 348 Index of advertisers

Supplements Computers & Typesetting: Errata and Changes, Supplement TUG Resource Directory

Volume 12, Number 2: Supplement ^{/} June 1991 TEX Users Group Resource Directory

3 TUG Membership List 118 TEX consulting and production services

TEX User Associations 119 General comments; TEX Users Group; CSTUG; CyrTUG; DANTE e.V.; GUTenberg; ITALIC; JTUG — TUG Japan; Nordic TEX Group; NTG; UK TUG; YUNUS – Turkish TEX users; Other international contacts

126 Elizabeth M. Barnhart ^{/} TEX L.U.G.s

Sources of TEX Software 127 Generic format; Acorn Archimedes; Amiga; Atari ST; Data General; DEC-20 — TOPS-20; IBM mainframes – MVS; IBM mainframes – VM/CMS; IBM PC and compatibles — MS-DOS; Macintosh; Prime; Unix; VAX/VMS

131 Ted Nieland ^{/} Announcing the DECUS TEX Collection, February 1991

Output Devices 133 Don Hosek ^{/} TEX Output Devices

TEX Archives 159 General observations; Aston (uk.ac.tex); Clarkson (L^{A}TEX styles);

e-math.ams.com; Heidelberg (DHDURZ1.Bitnet);
june.cs.washington.edu; labrea.stanford.edu;
Nijmegen (HEARN.Bitnet); SHSU (niord.shsu.edu , SHSU.bitnet);

SPAN/DECnet (Italy); Stuttgart (rusmv1.rus.uni-stuttgart.de);

Utrecht (archive.cs.ruu.nl); ymir.claremont.edu;

Other archives containing TEX material 164 R.M. Damerell ^{/} Fetching files from the Aston archive

Electronic Discussion Lists 170 General comments; comp.text.tex; ELLHNIKA (Greek); GUT (GUTenberg); INFO-TeX; IVRITEX (Semitic languages);

LaTeX-help; RUSTeX-L (Russian); TeX-D-L (German); TeX-EURO;
TeXhax; TEXMaG; TeX-Pubs; UKTeX; TeX-NL (Dutch);

YUNUS (Turkish)

TEX Bibliography 176 Publications about TEX and typography 183 Publications prepared with TEX

Volume 12, Number 3 &4 ^{/} November 1991 1991 TUG Conference Proceedings

Part 1

Introduction 351 Nelson H. F. Beebe ^{/} President’s introduction Keynote Address 353 Nico A.F.M. Poppelier ^{/} Two sides of the fence

Publishing 359 Laurie J. Petrycki ^{/} Comparing TEX and traditional typesetting for the composition of

a textbook

367 Frederick H. Bartlett ^{/} Contra-L^{A}TEX, or what really works in the publishing world 372 Howard Ratner and Kenneth Dreyhaupt ^{/} TEX in a book production department 377 Berthold K.P. Horn ^{/} dvi and eps: The ideal author-to-publisher interface? 382 Neil A. Weiss ^{/} Producing a book using TEX: How the process works 387 Samuel E. Rhoads ^{/} Authors new to TEX publish a textbook with a publisher new to

TEX

393 Colleen Brosnan ^{/} The “Five Cs”: A guide to successful publication using TEX 397 Anita Z. Hoover ^{/} L^{A}TEX/TEX user: A typist, or typesetter? 401 John Lavagnino ^{/} Simultaneous electronic and paper publication 406 Robert W. McGaffey ^{/} SGML versus/and TEX 409 Andrew E. Dobrowolski ^{/} Typesetting SGML documents using TEX 415 C. M. Sperberg-McQueen ^{/} Specifying document structure: Differences in L^{A}TEX and

TEI markup

422 Takashi Kakiuchi, Yuki Kusumi, Yoshiyuki Miyabe, and Kazu Tsuga ^{/} A structured document preparation system — Autolayouter version 2.0 — an enhancement for handling multiple document types

430 Linda Williams ^{/} Refining a process

Future Issues 434 Luigi Semenzato and Edward Wang ^{/} A text processing language should be first a

programming language

442 Michael Vulis ^{/} Should TEX be extended?

TUG Business 448 TUG 1992 annual meeting, Portland, Oregon

449 Institutional members

Forms 451 TUG membership applications

Advertisements 453 TEX consultants 468 Index of advertisers

Part 2

Applications 471 Cay S. Horstmann ^{/} Automatic conversion from a scientific word processor to TEX

479 Dennis S. Arnon and Sandra A. Mamrak ^{/} On the logical structure of mathematical

notation

485 Kees van der Laan ^{/} Math into BLUes 502 Michael J. Downes ^{/} Dialog with TEX 510 Jackie Damrau and Michael Wester ^{/} Form letters with 3-across labels capability 517 Mark A. Roth ^{/} Typesetting forms with L^{A}TEX 523 Peter Flynn ^{/} Developing a pop-up help facility for TEX on PCs 528 Malcolm Clark, Brian Hamilton Kelly, and Niel Kempson ^{/} 7 bits good, 8 bits bad or

“The eight-bit blight”

Graphics and Fonts 534 Friedhelm Sowa ^{/} Bitmaps and halftones with BM2FONT 539 Yannis Haralambous ^{/} TEX and those other languages 549 Don Hosek ^{/} Siamese TEX: Joining dvi files at the hip and other novel applications of

vf files

554 Alan Hoenig ^{/} When TEX and talk: Typesetting on curved paths and other special effects

Announcements 558 Participants at TUG’91 562 Calendar 563 The Donald E. Knuth Scholarship for 1992 564 GUTenberg’92, Les Diablarets, Switzerland 570 TUG 1992 annual meeting, Portland, Oregon

Advertisements 588 Index of advertisers

Volume 13, Number 1 ^{/} April 1992

3 Addresses

General Delivery 5 Prez says ^{/} Malcolm Clark 6 President’s introduction ^{/} Nelson H. F. Beebe 10 Editorial comments ^{/} Barbara Beeton 11 Samuel B. Whidden, 1930–1991

Software 13 Inside Type & Set ^{/} Graham Asher Philology 23 Computer Aided Hyphenation for Italian and Modern Latin ^{/} Claudio Beccari

Fonts 34 Invisibility using virtual fonts ^{/} Sebastian Rahtz 36 Packing s into PostScript ^{/} Toby Thain 39 Modern Greek with adjunct fonts ^{/} C. Mylonas and R. Whitney 51 Comments on “Filenames for Fonts” (TUGboat 11#4) ^{/} Frank Mittelbach

Output Devices 54 DVI driver standard, level 0 ^{/} TUG DVI Driver Standards Committee

Resources 57 New books on TEX ^{/} Victor Eijkhout

58 New books on L^{A}TEX ^{/} Nico Poppelier

Questions 60 Just plain Q&A: Of partitioned matrices and doublespacing ^{/} Alan Hoenig Tutorials 62 Elementary text processing and parsing in TEX — the appreciation of tokens — ^{/}

L. Siebenmann

Macros 74 The bag of tricks ^{/} Victor Eijkhout 75 Erratum: Oral TEX, TUGboat 12, no. 2, pp. 272–276 ^{/} Victor Eijkhout 75 Some basic control macros for TEX ^{/} Jonathan Fine 84 Self-replicating macros ^{/} Victor Eijkhout and Ron Sommeling 85 A font and a style for typesetting chess using L^{A}TEX or TEX ^{/} Piet Tutelaers 91 Tower of Hanoi, revisited ^{/} Kees van der Laan

L^{A}TEX 94 The L^{A}TEX column ^{/} Jackie Damrau

95 Erratum: “See also” indexing with Makeindex, TUGboat 12, no. 2, p. 290 ^{/}

Harold Thimbleby

96 L^{A}TEX 2.09 ↪ L^{A}TEX3 ^{/} Frank Mittelbach and Chris Rowley

Abstracts 101 Cahiers GUTenberg #9 and #10–11 News & Announcements 106 Calendar

107 EuroTEX 92, Prague, 14–18 September 1992

Late-Breaking News 108 Production notes ^{/} Barbara Beeton

109 Coming next issue

TUG Business 110 Institutional members

Forms 112 TUG membership application

Advertisements 126 Index of advertisers 127 TEX consulting and production services

TUGBOAT

Volume 13, Number 2 ^{/} July 1992

131 Addresses

General Delivery 133 Changing TEX? ^{/} Malcolm Clark 134 Editorial comments ^{/} Barbara Beeton 137 TUG seeks Executive Director 138 TEX: The next generation ^{/} Philip Taylor

Software 139 Knuth’s profiler adapted to the VMS operating system ^{/} R.M. Damerell

Fonts 146 Arrows for Technical Drawings ^{/} David Salomon Graphics 150 A solution to the color separation problem ^{/} Daniel Levin

156 A style option for rotated objects in TEX ^{/} Sebastian Rahtz and Leonor Barroca

Resources 181 Book review: An Italian guide to L^{A}TEX (by Claudio Beccari) ^{/}

Marisa Luvisetto and Massimo Calvani

182 Book reviews: Jane Hahn, L^{A}TEX for Everyone; Eric van Herwijnen,

Practical SGML ^{/} Nico Poppelier

185 Book review: Victor Eijkhout, TEX by Topic ^{/} Philip Taylor 188 A TEX macro index ^{/} David M. Jones

Tutorial 189 Names of control sequences ^{/} Victor Eijkhout Puzzle 190 Where does this character come from? ^{/} Frank Mittelbach Macros 191 The bag of tricks ^{/} Victor Eijkhout

192 Over the multi-column ^{/} Péter Huszár 201 The elementary Particle Entity Notation (PEN) scheme ^{/}

Michel Goossens and Eric van Herwijnen

L^{A}TEX 208 From TEX to L^{A}TEX ^{/} Maria Luisa Luvisetto and Enzo Ugolini 215 Geometric diagrams in L^{A}TEX ^{/} Peter J. Cameron 217 How to change the layout with L^{A}TEX 2.09 ^{/} Hubert Partl

SGML 221 SGML — Questions and answers ^{/} Reinhard Wonneberger and Frank Mittelbach Dreamboat 223 TEX wish list ^{/} Michael Barr

226 Approaching SGML from TEX ^{/} Reinhard Wonneberger

Abstracts 227 Cahiers GUTenberg #12

228 Baskerville, Volume 2, Number 1, March 1992

Late-Breaking News 229 Production notes ^{/} Barbara Beeton

230 Coming next issue

News & Announcements 231 Calendar

232 EuroTEX 92, Prague, 14–18 September 1992

Forms 235 TUG membership application

Advertisements 247 Index of advertisers Supplements TUG Membership List

Volume 13, Number 3 ^{/} October 1992 1992 Annual Meeting Proceedings

Introduction 251 Malcolm Clark ^{/} President’s introduction Keynote Address 253 Malcolm Clark ^{/} Portable graphics in TEX

Software 261 Bart Childs ^{/} Literate programming, a practitioner’s view

269 Steve Hampson and Barry Smith ^{/} A high performance TEX for the

Motorola 68000 processor family

Front Ends 272 Harry L. Baldwin, Jr. ^{/} Using a high-level language as an aid in writing TEX

documents

281 Larry F. Bennett ^{/} T-EDIT, a collection of editing macros for TEX 291 Robert McGaffey ^{/} Automatic tables using SGML, C, and TEX 295 Anthony J. Starks ^{/} Dotex — integrating TEX into the X-window system 304 Kresten Krab Thorup ^{/} GNU emacs as a front end to L^{A}TEX 309 Walter van der Laan and Johannes Braams ^{/} Writing reports with more than a

hundred people

Graphics 315 Jackie Damrau ^{/} Discovering graphics in L^{A}TEX documents 322 Robert L. Harris ^{/} Preparing halftones for use in TEX 327 David Salomon ^{/} Creating shaded rectangles with PostScript 330 Neil A. Weiss ^{/} Creation and incorporation of PostScript graphics with

TEX-formatted labels into TEX documents

Macros 335 Timo Knuutila ^{/} How to combine multiple languages, PostScript, and L^{A}TEX 341 Victor Eijkhout ^{/} Just give me a lollipop (it makes my heart go giddy-up) 347 James L. Hafner ^{/} FoilTEX, a L^{A}TEX-like system for typesetting foils 357 Peter Abbott ^{/} Typesetting a magazine the easy way

File Management 362 Mimi Burbank and Donna Burnette ^{/} Using TEX for a publications database

Future Issues 372 T. V. Raman ^{/} An audio view of (L^{A})TEX documents

380 Dennis S. Arnon, Isabelle Attali and Paul Franchi-Zannettacci ^{/}

Model-based conversions of L^{A}TEX documents

Reports 390 Chris Rowley ^{/} L^{A}TEX3 update

391 Workshops

Participants 393 Participants at the 1992 TUG Meeting

Announcements 395 The Donald E. Knuth Scholarship for 1992 and 1993 396 Calendar 398 TUG 1993 annual meeting, Aston University, UK 399 TUG 1993 course schedule

TUG Business 400 Institutional members Advertisements 402 Consultants

411 Index of advertisers

Volume 13, Number 4 ^{/} December 1992

415 Addresses

General Delivery 417 The back benches beckon ^{/} Malcolm Clark 418 Editorial comments ^{/} Barbara Beeton 419 An interview with Donald Knuth, November 1991

Dreamboat 425 Moving a fixed point ^{/} Richard Palais 433 The future of TEX ^{/} Philip Taylor

Software 443 XBibTEX and friends ^{/} Nickolas J. Kelly and Christian H. Bischof 447 Searching in a DVI file ^{/} Nigel Chapman 452 Hyphenation exception log

Literate programming 457 Errata: Literate Programming, A Practitioner’s View, TUGboat 13, no. 3, pp. 261–268 / Bart Childs

Philology 457 Hyphenation patterns for ancient Greek and Latin ^{/} Yannis Haralambous

Fonts 470 The exotic Croatian Glagolitic alphabet ^{/} Darko Žubrinić 472 Postnet codes using ^{/} John Sauter 476 A typewriter font for the Macintosh 8-bit font table ^{/} Yannis Haralambous

Graphics 477 Addendum: A style option for rotated objects in TEX (TUGboat 13, no. 2,

pp. 156–180) ^{/} Sebastian Rahtz and Leonor Barroca

478 Diag: a drawing preprocessor for L^{A}TEX ^{/} Ray Seyfarth

Book Reviews 486 Wynter Snow, TEX for the Beginner ^{/} Victor Eijkhout 487 Arvind Borde, TEX by Example ^{/} George Greenwade 489 André Heck, ed., Desktop Publishing in Astronomy & Space Sciences ^{/}

A. G. W. Cameron

Typesetting on PCs 490 TEX implementations for IBM PCs: comparative timings ^{/} Erich Neuwirth

Warnings 493 Where does this character come from? Solution to the puzzle, TUGboat 13, no. 2, p.190 ^{/} Frank Mittelbach

Macros 494 The bag of tricks ^{/} Victor Eijkhout 495 Too many errors ^{/} Jonathan Fine 496 One error less ^{/} Victor Eijkhout 497 ZzTEX: A macro package for books ^{/} Paul Anagnostopoulos 505 The \noname macros — A technical report ^{/} Jonathan Fine

L^{A}TEX 510 Volunteer work for the L^{A}TEX3 project ^{/} Frank Mittelbach,

Chris Rowley and Michael Downes

516 Correction sheets in L^{A}TEX ^{/} Mike Piff 518 Text merges in TEX and L^{A}TEX ^{/} Mike Piff 524 A style file for printing sheets of labels ^{/} Sebastian Rahtz

Abstracts 528 Cahiers GUTenberg #13

News & Announcements 530 Calendar

532 Call for Papers: Special Issue of Electronic Publishing: Origination, Dissemination and Design on Active Documents

Late-Breaking News 533 Production notes ^{/} Barbara Beeton 534 Coming next issue

TUG Business 534 Institutional members

Forms 537 TUG membership application

Advertisements 536 Index of advertisers 544 TEX consulting and production services

TUGBOAT

Volume 14, Number 1 ^{/} April 1993

3 Addresses

General Delivery 5 Opening words ^{/} Christina Thiele

7 Editorial comments ^{/} Barbara Beeton

Philology 8 Typesetting Chinese pinyin using virtual fonts ^{/} Wai Wong

Hardware / Systems 12 A multimedia document system based on TEX and DVI documents ^{/}

R. A. Vesilo and A. Dunn

Book Reviews 17 Arvind Borde, Mathematical TEX by Example ^{/} Philip Taylor

20 Michael Vulis, Modern TEX and Its Applications ^{/} Jon Radel

Macros 23 The bag of tricks ^{/} Victor Eijkhout 25 Anchored figures at either margin ^{/} Daniel Comenetz 35 The \CASE and \FIND macros ^{/} Jonathan Fine 40 Doing astronomical computations with TEX: Making agendas ^{/} Jordi Saludes 54 FIFO and LIFO sing the BLUes ^{/} Kees van der Laan

L^{A}TEX 60 An update on the babel system ^{/} Johannes Braams

62 Hacker’s Guide to AMSFonts and NFSS in the Context of L^{A}TEX ^{/}

Rafał Żbikowski

Letters 70 Response to A.G.W. Cameron ^{/} André Heck Abstracts 71 Die TEXnische Komödie 1992, Heft 1–4

News & Announcements 77 Calendar

79 Courses to be held in conjunction with TUG 93

(Aston University, Birmingham, U.K., 26–30 July 1993)

Late-Breaking News 81 Production notes ^{/} Barbara Beeton

81 Coming next issue

TUG Business 83 Institutional members

Forms 85 TUG membership application

Advertisements 82 Index of advertisers 87 TEX consulting and production services

TUGBOAT

Volume 14, Number 2 ^{/} July 1993

95 Addresses

General Delivery 97 Opening words ^{/} Christina Thiele

99 Editorial comments ^{/} Barbara Beeton TEX 3.1415 / 2.71 on the way; L^{A}TEX 2e coming in October; Annual meeting at Aston University; CTAN — the Comprehensive TEX Archive Network; TUGboat tables of contents on-line; Encoding of 256-character math fonts; Comment on font naming

102 The Donald E. Knuth Scholarship: 1993 Scholar and 1994 announcement ^{/}

Nico Poppelier

103 The A-in-L^{A}TEX Contest: Deadline extended

Philology 104 Fonts for Africa: the fc-fonts ^{/} Jörg Knappen

Fonts 107 Implementing the extended TEX layout using PS fonts ^{/} Sebastian Rahtz

118 Zebrackets: a pseudo-dynamic contextually adaptive font ^{/} Michael Cohen

Graphics 123 From observation to publication ^{/} Theo Jurriens Book Reviews 127 Review of recent L^{A}TEX books ^{/} Nico Poppelier

Helmut Kopka and Patrick Daly, A Guide to L^{A}TEX; Antoni Diller, L^{A}TEX Line By Line

Typesetting on PCs 128 ET — a TEX-compatible editor for MSDOS computers ^{/} John Collins

Tutorials 132 Essential NFSS2 ^{/} Sebastian Rahtz

138 A pragmatic approach to paragraphs ^{/} Philip Taylor

Letters 141 Truth in indexing ^{/} Lincoln Durst Macros 141 Letterspacing in TEX ^{/} Philip Taylor

Abstracts 146 Cahiers GUTenberg #15

News & Announcements 147 Calendar

Late-Breaking News 149 Production notes ^{/} Barbara Beeton

149 Coming next issue

TUG Business 150 Institutional members Advertisements 148 Index of advertisers

152 TEX consulting and production services

Volume 14, Number 3 ^{/} October 1993 1993 Annual Meeting Proceedings

Keynote Addresses 162 Christina A.L. Thiele ^{/} The future of TEX and TUG 167 Joachim Lammarsch ^{/} A new typesetting system: is it really necessary? 171 Marek Ryćko and Bogusław Jackowski ^{/} TEX from \indent to \par

Futures 177 Philip Taylor ^{/} NTS: the future of TEX? 183 Roger Hunter ^{/} A future for TEX

Macro Packages 187 Frank G. Bennett, Jr. ^{/} LexITEX: context-sensitive legal citations for L^{A}TEX 196 Daniel Taupin ^{/} Using TEX and to build complicated maps 203 Daniel Taupin ^{/} MusicTEX: Using TEX to write polyphonic or

instrumental music

212 Laurent Siebenmann ^{/} A format compilation framework for European languages

Working with TEX 222 Nelson Beebe ^{/} Bibliography prettyprinting and syntax checking

223 Martin Bryan ^{/} A TEX user’s guide to ISO’s Document Style Semantics and

Specification Language (DSSSL)

227 Peter Flynn ^{/} TEX and SGML: a recipe for disaster? 231 Xinxin Wang and Derick Wood ^{/} An abstract model for tables 238 Michel Lavaud ^{/} Developing a multi-windowing environment for research

based on TEX

Drivers 245 Minato Kawaguti ^{/} A versatile TEX device driver

Language Issues 252 Gabriel Valiente Feruglio and Robert Fuster ^{/} Typesetting Catalan texts

with TEX

260 Yannis Haralambous ^{/} The Khmer script tamed by the Lion (of TEX) 271 John Plaice ^{/} Language-dependent ligatures

Fonts 275 Michael Doob and Craig Platt ^{/} Virtual fonts in a production environment 282 Berthold K.P. Horn ^{/} Where are the math fonts? 285 Alan Jeffrey ^{/} A PostScript font installation package written in TEX 293 (edited by Alan Jeffrey) ^{/} Math font encodings: a workshop summary

Macro Writing 296 Włodek Bzyl and Tomasz Przechlewski ^{/} An application of literate

programming: creating a format for the Bulletin of the Polish TUG

300 Jonathan Fine ^{/} Galleys, space, and automata 310 Kees van der Laan ^{/} Syntactic sugar 319 Kees van der Laan ^{/} Sorting within TEX

Education and Archives 329 Mary Dyson ^{/} Teaching digital typography — the Didot Project

333 Irina V. Gorbunova ^{/} Russian-speaking user: from Chi-Writer and

Ventura Publisher to TEX; learning difficulties

335 R. Allan Reese ^{/} How to set up and maintain a TEX system 342 George D. Greenwade ^{/} The Comprehensive TEX Archive Network (CTAN)

Delegates 352 Participants at the 1993 TUG meeting

TUG Business 356 Institutional members

Announcements 358 Calendar 360 TUG 1994 annual meeting, Santa Barbara, USA

Forms 361 TUG individual membership application 362 TUG institutional membership application

Advertisements 359 Index of advertisers 363 Consultants

TUGBOAT

Volume 14, Number 4 ^{/} December 1993

369 Addresses

General Delivery 371 Opening words ^{/} Christina Thiele

TEX at meetings of other societies; Free-Net; Renew for 1994

372 Editorial comments ^{/} Barbara Beeton

Reminder to potential TUGboat authors; Call for volunteers

373 TUGboat wish list

Dreamboat 374 NEXTEX: A personal view ^{/} Malcolm Clark

381 NTS update ^{/} Philip Taylor

Software & Tools 382 Two extensions to GNU Emacs that are useful when editing TEX documents ^{/}

Thomas Becker

387 Icons for TEX and ^{/} Donald E. Knuth 390 bibview: A graphical user interface to BibTEX ^{/} Armin Liebl 395 Bibliography prettyprinting and syntax checking ^{/} Nelson Beebe

Graphics 420 A tough table becomes easy with PICTEX ^{/} Kevin Carmody Book Reviews 421 Book Review: P. W. Abrahams, K. Berry and K. Hargreaves,

TEX per l’impaziente ^{/} Claudio Beccari

Hints & Tricks 423 Ten TEX tricks for the mathematician ^{/} Helmer Aslaksen Macros 424 The bag of tricks ^{/} Victor Eijkhout L^{A}TEX 425 The “operational requirement” for support of bibliographies ^{/} David Rhead

433 Relative moves in L^{A}TEX pictures ^{/} Richard Bland

News & Announcements 438 Calendar

Late-Breaking News 439 Production notes ^{/} Barbara Beeton

440 Coming next issue

TUG Business 441 Institutional members

Forms 443 TUG membership application

Advertisements 442 Index of advertisers 445 TEX consulting and production services

Volume 15, Number 1 ^{/} March 1994

3 Addresses

General Delivery 5 Opening words ^{/} Christina Thiele Board news; The annual meeting; TEX in new places; Announcing EuroTEX’94; Office updates

6 Editorial comments ^{/} Barbara Beeton TUGboat wish list; Another award for DEK; New magazine on type and typography; Hyphenation and exceptions

7 The TEX hierarchy ^{/} Donald Arseneau, Raymond Chen and Victor Eijkhout

Philology 9 Typesetting of ancient languages ^{/} Claudio Beccari

17 Comments on the paper “Typesetting Catalan texts with TEX” (TUGboat 14,

no. 3, pp. 252–259) ^{/} Claudio Beccari

17 Comments on the comments: Typesetting Catalan texts with TEX ^{/} Gabriel Valiente Feruglio and Robert Fuster

Book Reviews 18 George Grätzer, Math into TEX ^{/} Nico Poppelier 19 Stephan von Bechtolsheim, TEX in Practice ^{/} T. L. (Frank) Pappas 21 Gianni Gilardi, Il TEX – Introduzione al linguaggio e complementi avanzati ^{/}

Claudio Beccari

22 Erik Spiekermann & E. M. Ginger, Stop Stealing Sheep ^{/} Merry Obrecht Sawdey 24 Eric van Herwijnen, Practical SGML ^{/} Nico Poppelier 25 Donald E. Knuth, Literate Programming ^{/} Christine Detig and Joachim Schrod

Tutorials 28 Output routines: Examples and techniques, Part IV: Horizontal techniques ^{/}

David Salomon

40 Verbatim copying and listing ^{/} David Salomon

Macros 55 The bag of tricks ^{/} Victor Eijkhout 57 Random bit generator in TEX ^{/} Hans van der Meer 59 Slanted lines with controlled thickness ^{/} David Salomon

and Matthew N. Hendryx

Letters 62 On indexing and errors ^{/} Jonathan Fine

L^{A}TEX 63 L^{A}TEX2ε update, dateline: 31 January 1994 ^{/} Chris Rowley

64 Addenda: A suggested “operational requirement” for L^{A}TEX3’s treatment of bibliographic references (TUGboat 14, no. 4, pp. 425-433) ^{/} David Rhead

News & Announcements 66 Calendar

68 TUG ’94 — Announcement and preliminary program

(Santa Barbara, California, 31 July–4 August 1994)

69 Call for papers, EuroTEX ’94 (Gdansk, Poland, 26–30 September 1994)

Late-Breaking News 74 Production notes ^{/} Barbara Beeton 75 Coming next issue

TUG Business 70 Meet the Board, Part I Barbara Beeton; Mimi Burbank; Jackie Damrau; Luzia Dietsche; Michael Doob; Michel Goossens; Tom Rokicki

75 Institutional members

Forms 77 TUG membership application

Advertisements 79 TEX consulting and production services 81 Index of advertisers

Volume 15, Number 2 ^{/} June 1994

87 Addresses

General Delivery 89 Opening words ^{/} Christina Thiele

Meetings; ISO and de facto standards

89 Editorial comments ^{/} Barbara Beeton “L^{A}TEX2ε” is now just “L^{A}TEX”; New CTAN features; Northwest Computing Support Center closed; TEX output in audible form

92 WEPT: A Week on Electronic Publishing and Typography ^{/} Michel Goossens

Dreamboat 96 Report of the 2nd meeting of the NTS group, February 1994 ^{/} Philip Taylor

Fonts 97 {Meta}Font Forum redux ^{/} Alan Hoenig

98 ASCII.sty: A new style-option and encoded font with IBM graphics control characters for use with TEX and L^{A}TEX ^{/} R. Ramasubramanian, R.W.D. Nickalls and M.A. Reed

Book Reviews 103 Michel Goossens, Frank Mittelbach, and Alexander Samarin,

The L^{A}TEX Companion ^{/} David M. Jones and David E. Wald

106 Norman Walsh, Making TEX Work ^{/} Victor Eijkhout 107 Christian Rolland, L^{A}TEX guide pratique ^{/} Jacques André

Typesetting on Personal Computers 107 NextTEX: TEX plus the NextStep Operating System ^{/} Alan Hoenig

Macros 110 Interaction tools: dialog.sty and menus.sty ^{/} Michael Downes L^{A}TEX 131 L^{A}TEX V3: philology & typography, reports to read, reports to do ^{/}

Bernard Gaulle

131 Tough table becomes easy with PICTEX, but it’s even easier with L^{A}TEX ^{/}

Claudio Beccari

Letters 132 On the review of TEX in Practice ^{/} Paul Anagnostopolous Abstracts 133 Baskerville, Volume 4, Nos. 1–2

136 Cahiers GUTenberg Nos. 16 and 17 138 Die TEXnische Komödie 1993, Heft 1–4

News & Announcements 143 Calendar

160 TUG ’95 — St. Petersburg, Florida

Late-Breaking News 148 Production notes ^{/} Barbara Beeton 149 Coming next issue 150 Change in TUGboat policy ^{/} Barbara Beeton

TUG Business 145 Meet the Board, Part II Michael Ferguson; Peter Flynn; George Greenwade; Yannis Haralambous; Nico Poppelier; Jon Radel; Sebastian Rahtz

151 Institutional members

Forms 155 TUG membership application

Advertisements 152 TEX consulting and production services 154 Index of advertisers

Volume 15, Number 3 ^{/} September 1994 1994 Annual Meeting Proceedings

166 Editorial and production notes 167 Acknowledgements and Conference Program: Innovation

Keynote Publishing, Languages Literature and Fonts 169 Charles Bigelow ^{/} Lucida and TEX: lessons of logic and history

170 Frank Mittelbach and Michel Goossens ^{/} Real life book production — lessons learned

from The L^{A}TEX Companion

174 Yannis Haralambous ^{/} Typesetting the holy Bible in Hebrew, with TEX 192 Michel Cohen ^{/} Adaptive character generation and spatial expressiveness 199 Yannis Haralambous ^{/} Humanist 200 Basil Malyshev ^{/} Automatic conversion of fonts to Type1 PostScript

Colour, and L^{A}TEX 201 James Hafner ^{/} The (pre)history of color in Rokicki’s dvips 205 Tom Rokicki ^{/} Advanced ‘special’ support in a dvi driver 213 Angus Duggan ^{/} Colour separation and PostScript 218 Sebastian Rahtz and Michel Goossens ^{/} Simple colour design with L^{A}TEX2ε 223 Friedhelm Sowa ^{/} Printing colour pictures 228 Michael Sofka ^{/} Color book production using TEX 239 Timothy van Zandt and Denis Girou ^{/} Inside PSTricks 247 Jon Stenerson ^{/} A L^{A}TEX style file generator 255 Johannes Braams ^{/} Document classes and packages in L^{A}TEX2ε 263 Alan Jeffrey ^{/} PostScript font support in L^{A}TEX2ε

TEX Tools 269 Oren Patashnik ^{/} BibTEX 1.0 274 Pierre MacKay ^{/} A typesetter’s toolkit 285 Michael P. Barnett and Kevin R. Perry ^{/} Symbolic computation for electronic

publishing

293 Minato Kawaguti and Norio Kitajima ^{/} Concurrent use of interactive TEX previewer

with an Emacs-type Editor

301 Yannis Haralambous ^{/} An Indic TEX preprocessor — Sinhalese TEX 302 Jean-Luc Doumont ^{/} Pascal pretty-printing: an example of “preprocessing within

TEX”

Futures 309 Joachim Schrod ^{/} Towards interactivity for TEX 318 Chris Rowley and Frank Mittelbach ^{/} The floating world 319 Don Hosek ^{/} Sophisticated page layout with TEX 320 John Plaice ^{/} Progress in the Omega project 325 Arthur Ogawa ^{/} Object-oriented programming, descriptive markup, and TEX 331 William Erik Baxter ^{/} An object-oriented programming system in TEX 339 Norm Walsh ^{/} A World Wide Web interface to CTAN 344 Yannis Haralambous and John Plaice ^{/} First applications of Ω: Adobe Poetica,

Arabic, Greek, Khmer

353 Philip Taylor ^{/} ε-TEX & NT S: A progress report

Publishing and Design 359 Maurice Laugier and Yannis Haralambous ^{/} TEX innovations by the Louis-Jean

Printing House

360 Michael Downes ^{/} Design by template in a production macro package 369 Alan Hoenig ^{/} Less is More: Restricting TEX’s scope enables complex page layouts 381 Jonathan Fine ^{/} Documents, compuscripts, programs and macros 386 Marko Grobelnik, Dunja Mladeniç, Darko Zupaniç and Borut Žnidar ^{/}

Integrated system for encyclopaedia typesetting based on TEX

388 Henry Baragar and Gail E. Harris ^{/} An example of a special purpose input language to L^{A}TEX

Appendix 397 Color pages 405 Participants at the Annual Meeting

News & Announcements 408 Calendar 409 1995 Knuth Scholarship

TUG Business 411 Institutional members 415 TUG membership application

Advertisements 412 TEX consulting and production services 417 Index of advertisers

TUGBOAT

Volume 15, Number 4 ^{/} December 1994

423 Addresses

General Delivery 425 Opening words ^{/} Christina Thiele

426 Editorial comments ^{/} Barbara Beeton TEX meetings in 1994; A new, expanded TEX FAQ; UKTeX Digest ceases to exist; Miscellaneous gossip

427 The TUG94 Proceedings — Apologia ^{/} Michel Goossens,

Sebastian Rahz and Barbara Beeton

429 EuroTEX94 ^{/} Julita Bolland, Toni Walter and Włodek Bzyl

Dreamboat 434 Minutes of the NTS meeting held at Lindau on October 11/12th 1994 ^{/}

Philip Taylor

Tools 438 TEX innovations at the Louis-Jean printing house ^{/} Maurice Laugier

and Yannis Haralambous

Fonts 444 Recycled ^{/} Pierre A. MacKay

447 Indica, an Indic preprocessor for TEX: A Sinhalese TEX system ^{/}

Yannis Haralambous

Graphics 459 The EAN barcodes by TEX ^{/} Petr Olšák Letters 464 An open letter to the TUG Board ^{/} Philip Taylor Survey 466 Typesetting commutative diagrams ^{/} Gabriel Valiente Feruglio Macros 485 The bag of tricks ^{/} Victor Eijkhout

486 A TEX autostereogram generator ^{/} Jacques Richer 492 Stereographic pictures using TEX ^{/} Reinhard Fößmeier

L^{A}TEX 496 To reset or not to reset ^{/} Johannes Braams Abstracts 498 Cahiers GUTenberg Nos. 18 and 19

500 Abstracts of the Proceedings of the Eighth European TEX Conference,

Gdańsk, September 1994

News & Announcements 505 Calendar

505 Production notes ^{/} Barbara Beeton

TUG Business 506 Institutional members

Forms 507 TUG membership application

Advertisements 508 TEX consulting and production services